OpenStudy (anonymous):

452.7 x 10^3 - 2.7 x 10^5

OpenStudy (mathmale):

Remember order of operations rules? As they apply here, you must perform exponentiation before you multiply, and multiply before you subtract. Give this a try.

OpenStudy (phebe):

Its 182700 but @mathmale will show u how to get it XD

TheSmartOne (thesmartone):

452.7 x10^3 =452700 2.7 x10^5 =270000 So 452700-270000=182700

TheSmartOne (thesmartone):

@hunter_225

OpenStudy (mathmale):

@hunter_225 : Need to hear from you. What is 10^3? 10^5?

OpenStudy (phebe):

@TheSmartOne thas what i got too

OpenStudy (mathmale):

@hunter_225 : Please show some signs of life if you still want help here.